OpenStudy (anonymous):

Help: Show all steps. Solve the system by elimination. -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5

OpenStudy (anonymous):

@wio Canyou help me with this please?

OpenStudy (campbell_st):

adding the 1st and 3rd equations then 2nd and 3rd equations will give 2 equations in 2 unknowns and eliminate x 1 + 3 5y + 6z = 5 equations 1 + 3 call it equation 4 2y + 4z = 2 Equations 2 + 3 call the result equation 5 double equation 4 and triple equation 5 and then subtract them to eliminate z 10y + 12z = 10 - 6y +12z = 6 -------------------- 4y = 4 I'll leave you to find y when you have it substitute into equation 4 to find z... when you have y and z substitute into equation 1 to find x... hope it helps

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Would y=1? @campbell_st

OpenStudy (campbell_st):

thats correct...

OpenStudy (anonymous):

@campbell_st can you help me with another one?

OpenStudy (campbell_st):

if you post it in the normal manner I'll look at it, and there are plenty of others who can help

OpenStudy (anonymous):

I posted it as another question.